year, x tuition, y The data in the chart show the cost of tuition at a state university for the years 1997-2001. Fit a regression line to the data, where x is the number of years after 1997. 1. (1998) 10,287 Then find the coefficient of correlation. Finally, use the 2. (1999) 10,787 0. (1997) 9651 Choose the correct regression line. O A. y-610.7x+9621 O C. y 610.7x-9621 O B. y 9621x+610.7 O D. y 9621 The coefficient of correlation is (Round to the nearest thousandth as needed) Click to select your answer(s)
university for the years 1997-2001. Fit a regression line to the data, where x is the number of years after 1997. 1 (1998) 10,287 Then find the coefficient of correlation. Finally, use the 2. (1999) 10,787 regression line to predict the cost of tuition in 2004 2007, and 2010 3 (2000) 11,278 4. (2001) 12,209 Is the regression line a good fit? O No O Yes The predicted cost of tuition in 2004 is $ (Round to the nearest dollar as needed.)
The predicted cost of tuition in 2007 is $ Round to the nearest dollar as needed.) The predicted cost of tuition in 2010 is S . (Round to the nearest dollar as needed.)
Answer #1

a) The correct regression line is : y = 610.7x+9621

Since this line gives values of y for x = 0, 1, 2, 3, 4 which are closest than the other lines.

b) COMPUTATION TABLE :

x y x2 xy y2
0 9651 0 0 93141801
1 10287 1 10287 105822369
2 10787 4 21574 116359369
3 11278 9 33834 127193284
4 12209 16 48836 149059681
10 54212 30 114531 591576504

Now, the coefficient of correlation is :

n(Σ.ry) _ (Σ.r ) ( Σ y)

i.e., 5 114531 1054212 5 *30-10]5 * 591576504 - 542122

i.e., 30535 50 18941576)

i.e., 30535 V947078800

i.e., ะ 0.992

Therefore, the coefficient of correlation is = 0.992.

c) Since the coefficient of correlation is very close of 1, therefore the regression line is a good fit.

d) Putting x = 7 in the above equation we get,

y(7) = 610.7*7+9621

i.e., y(7) approx 13896

Therefore, the predicted cost of tuition in 2004 is $13896.

e) Putting x = 10 in the above equation we get,

y(7) = 610.7*10+9621

i.e., y(7) = 15728

Therefore, the predicted cost of tuition in 2004 is $15728.

f) Putting x = 13 in the above equation we get,

y(7) = 610.7*13+9621

i.e., y(7) approx 17560

Therefore, the predicted cost of tuition in 2004 is $17560.